OpenStudy (anonymous):

can anyone help me please in this precal question? write the polynomial in standard from 3,-13,and 5+4i.. please and thank you

OpenStudy (anonymous):

start with \[(x-3)(x+13)\] since that will give you the zeros of \(3\) and \(-13\)

OpenStudy (anonymous):

then you have to multiply that by the quadratic that has zeros \(5+4i\) and \(5-4i\) that one is not too hard to find do you know how to do it?

OpenStudy (anonymous):

i can show you if you like two ways, one is easy the other is real real easy

OpenStudy (anonymous):

yea can you show me the real easy way?

OpenStudy (anonymous):

ok but first the easy way, because the real real easy way requires memorizing something

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

start with \[x=5+4i\] and work backwards subtract \(5\) and get \[x-5=4i\] then square both sides (carefully) get \[(x-5)^2=(4i)^2\\ x^2-10x+25=-16\] then add \(16\) and get \[x^2-10x+41\] as your quadratic

OpenStudy (anonymous):

that is the easy enough way, like solving a quadratic equation only backwards the real real easy way (if you are going to do more than one of these) is to memorize that if \[a+bi\] is a zero then the quadratic is \[x^2-2ax+(a^2+b^2)\]

OpenStudy (anonymous):

you had \[5+4i\] so the quadratic was \[x^2-2\times 5x+(5^2+4^2)=x^2-10x+41\]

OpenStudy (anonymous):

final job is to multiply out \[(x-3)(x+13)(x^2-10x+41)\] and again although this is annoying there is a real real easy way to do iot

OpenStudy (anonymous):

ok lemme try workin it out sorry all new to this stuff lol

OpenStudy (anonymous):

ok thank you so much! I got it now

OpenStudy (anonymous):

yw